Approximately what size wings would an aircraft need on Mars if its engine generates the same differences in speed as in the "Practice It" ( wing area: 3.50 m2, air flows over the top of the wing at 246 m/s, air flows underneath the wing at 223 m/s) and the total mass of the craft is 300 kg? The density of air on the surface of Mars is approximately one percent Earth's density at sea level, and the acceleration of gravity on the surface of Mars is about 3.8 m/s2.
Answer #1

apply the formula Pressure P = Force/Area = 0.5 rho (V2^2 - v1^2)

P = 0.5* 1200*1/100 * (246^2 - 223^2)

P = F/A = 1380

Area = F/6472200

A = mg/6472200

A = 300 * 3.8/(1380)

A =0.826 m^2